OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5777 — Optical pendulum migration Haubourdin (HAU59)
Scheduled Maintenance Report for Web Cloud
Completed
Hi,

We switch the site of Haubourdin on the new optical infrastructure in the night from Tuesday to Wednesday. Work will begin Tuesday evening at 0:00. The interruption for customers located on this site is estimated at 1:00. We will regularly update the task to keep you informed about the progress of work.


Update(s):

Date: 2011-09-23 11:08:41 UTC
The connections are operational. HAU59 is UP again.

Date: 2011-09-23 11:07:57 UTC
We have just adjusted the optical target between roubaix and Haubourdin to achieve the operating range of the equipment. The light goes well between Haubourdin and roubaix

Date: 2011-09-23 10:51:05 UTC
We will turn off the haubourdin site to switch on the optical equipment.The connexions will be disrupted during the intervention.

Date: 2011-09-23 10:40:44 UTC
We will start the intervention shortly.

Date: 2011-09-23 10:40:15 UTC
We will turn on the site of Haubourdin this night at 0:00

Date: 2011-09-20 05:21:17 UTC
We encountered a problem with the optics. After several investigations, we have not solved the problem.Intervention reported.

Date: 2011-09-19 18:37:50 UTC
We will switch the site of haubourdin tonight. the working will
start at 1:00 am.

Date: 2011-09-19 18:36:54 UTC
The intervention is reported to next week. We will give you later
the exact date of the migration in this working task.
Posted Sep 12, 2011 - 13:59 UTC